OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope point: (-3,7); slope:4

OpenStudy (anonymous):

choices?

OpenStudy (akashdeepdeb):

Use the point slope form of a line: \(y-y_1 = m(x-x_1)\) Where the point is = \((x_1,y_1)\) Where the slope is = m Understood? :)

OpenStudy (anonymous):

I still don't get it

OpenStudy (akashdeepdeb):



OpenStudy (anonymous):

Y-7=4(x-3) Y-7=4(x+3) Y+7=4(x+3) Y+7=4x-12

OpenStudy (anonymous):

Please help

OpenStudy (akashdeepdeb):

Plug in the values like I said in the first comment. CHECK.